OpenStudy (anonymous):

use the graph of the integrand and areas to evaluate the integral: upper limit 2, lower limit -2, f(x)=(5-lxl). can you explain how to evaluate this intergral?

OpenStudy (anonymous):

\[\int\limits_{-2}^{2}(5-\left| x \right|) dx\]

OpenStudy (anonymous):

Evaluate as 2 integrals. \[\int\limits_{-2}^{0}(x+5)dx + \int\limits_{0}^{2}(5-x)dx\]

OpenStudy (anonymous):

My problem is that I've never learned that as well. I have gone as far as learning to do it with a single number there, (the x is not there), kind of questions. We were jumped into these problems and harder ones

OpenStudy (anonymous):

You would increase the power of x by 1 and divide by that value, making x^2/2. 5 becomes 5x. Evaluate x^2/2 + 5x from -2 to 0 for the first integral.

OpenStudy (anonymous):

So.. it would then become ie. (-1^2/2) +5(-1) and then i just solve? And from there I do the same with the other side?